| #193 |
13 Apr 14:23 Log created
|
ryanque
|
Start experiment
|
normal
|
|
|
| #194 |
13 Apr 14:25 Antenna VSWR
|
ryanque
|
Antennas
|
normal
|
|
|
| #195 |
14 Apr 07:15 Restart of the experiment
|
I. Manay
|
Start experiment
|
normal
|
|
|
| #196 |
14 Apr 07:43 Fall of the transmitters
|
I. Manay
|
Transmitters
|
normal
|
|
|
| #197 |
14 Apr 09:55 adjustment of the vipir clock
|
I. Manay
|
General
|
normal
|
|
|
| #198 |
14 Apr 14:42 Interference at channel A
|
I. Manay
|
Reception
|
normal
|
|
|
| #199 |
14 Apr 18:36 Spark-gaps summary
|
W. Rojas
|
Spark Gaps
|
normal
|
|
|
| #200 |
14 Apr 19:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #201 |
14 Apr 19:59 Operation resume
|
I. Manay
|
General
|
normal
|
|
|
| #202 |
14 Apr 23:05 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #203 |
14 Abr 20:00 Ch B had high noise
|
P. Condor
|
Acquisition system
|
normal
|
|
|
| #204 |
14 Abr 23:59 Plots of rtis and spectras.
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #205 |
15 Abr 05:00 Spark gaps adjustment
|
A. Huamani
|
Spark Gaps
|
normal
|
|
|
| #206 |
15 Apr 07:05 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #207 |
15 Abr 08:00 Operation Resume
|
P. Condor
|
General
|
normal
|
|
|
| #209 |
15 Abr 10:56 Transmitter turned off.
|
|
Transmitters
|
normal
|
|
|
| #210 |
15 Abr 11:38 Transmitters turned on again.
|
|
Transmitters
|
normal
|
|
|
| #211 |
15 Abr 11:38 Spark gaps adjustment
|
|
Spark Gaps
|
normal
|
|
|
| #212 |
15 Abr 11:40 Channel 4 cable adjustment.
|
|
Reception
|
normal
|
|
|
| #213 |
15 Abr 13:40 Acquisition software halted
|
|
Acquisition system
|
normal
|
|
|
| #214 |
15 Abr 14:00 Acquisition software halted
|
|
Acquisition system
|
normal
|
|
|
| #215 |
15 Abr 14:22 Spark gaps adjustmen.
|
|
Spark Gaps
|
normal
|
|
|
| #216 |
15 Abr 16:00 Acquisition system hard drive changed.
|
|
Acquisition system
|
normal
|
|
|
| #217 |
15 Abr 16:10 End of the hard drive change.
|
|
Acquisition system
|
normal
|
|
|
| #218 |
15 Apr 19:50 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #219 |
15 Abr 19:30 Spark gaps adjustment.
|
|
Spark Gaps
|
normal
|
|
|
| #220 |
15 Abr 19:50 Operation resume
|
|
General
|
normal
|
|
|
| #221 |
15 Apr 23:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #222 |
15 Abr 23:59 RTI and Spectra Plots
|
|
Plot
|
normal
|
|
|
| #223 |
16 Apr 07:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #224 |
16 Apr 06:15 Spark gaps summary
|
|
Spark Gaps
|
normal
|
|
|
| #225 |
16 Abr 08:00 Operation Resume
|
|
General
|
normal
|
|
|
| #226 |
16 Apr 15:14 Spark gap summary
|
ryanque
|
Antennas
|
normal
|
|
|
| #227 |
16 Apr 15:25 Transmitter failure
|
ryanque
|
Transmitters
|
normal
|
|
|
| #228 |
16 Apr 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #229 |
16 Apr 20:15 Transmitter failures
|
ryanque
|
Transmitters
|
normal
|
|
|
| #230 |
16 Apr 20:22 Operation Resume
|
ryanque
|
General
|
normal
|
|
|
| #231 |
16 Abr 21:27 Problem with transmitter
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #232 |
17 Abr 00:28 Problem with transmitter
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #233 |
17 Abr 00:57 Transmitter is on
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #234 |
17 Abr 00:00 Plots of rtis and spectras
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #235 |
17 Apr 01:54 Problem with transmitter
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #236 |
17 Apr 02:16 Transmitter is on.
|
P. Condor
|
Transmitters
|
normal
|
|
|
| #237 |
17 Apr 07:00 Spark gaps adjustment
|
A. Coronado
|
Spark Gaps
|
normal
|
|
|
| #238 |
17 Apr 07:59 Operation Resume
|
P. Condor
|
General
|
normal
|
|
|
| #239 |
17 Apr 07:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #240 |
17 Apr 12:46 Fall of the transmitters
|
I. Manay
|
Transmitters
|
normal
|
|
|
| #241 |
17 Apr 18:36 Spark-gaps summary
|
I. Manay
|
Spark Gaps
|
normal
|
|
|
| #242 |
17 Apr 19:59 Operation resume
|
I. Manay
|
General
|
normal
|
|
|
| #243 |
17 Apr 20:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #244 |
05 May 11:40 Operativity report
|
|
General
|
normal
|
|
|
| #448 |
30 Jul 13:00 Resume MST-ISR-EEJ Apr 2015
|
|
End experiment
|
normal
|
|
|